# Autonomous Data Agents: A New Opportunity for Smart Data

### v1

@@ Abstract.
As data continues to grow in scale and complexity, preparing, transforming, and analyzing it remains labor-intensive, repetitive, and difficult to scale. Since data contains knowledge and AI learns knowledge from it, the alignment between AI and data is essential. However, data is often not structured in ways that are optimal for AI utilization. Moreover, an important question arises: how much knowledge can we pack into data through intensive data operations? Autonomous data agents (DataAgents), which integrate LLM reasoning with task decomposition, action reasoning and grounding, and tool calling, can autonomously interpret data task descriptions, decompose tasks into subtasks, reason over actions, ground actions into python code or tool calling, and execute operations. Unlike traditional data management and engineering tools, DataAgents dynamically plan workflows, call powerful tools, and adapt to diverse data tasks at scale. This report argues that DataAgents represent a paradigm shift toward autonomous data-to-knowledge systems. DataAgents are capable of handling collection, integration, preprocessing, selection, transformation, reweighing, augmentation, reprogramming, repairs, and retrieval. Through these capabilities, DataAgents transform complex and unstructured data into coherent and actionable knowledge. In this report, we first examine why the convergence of agentic AI and data-to-knowledge systems has emerged as a critical trend. We then define the concept of DataAgents and discuss their architectural design, training strategies, as well as the new skills and capabilities they enable. Finally, we call for concerted efforts to advance action workflow optimization, establish open datasets and benchmark ecosystems, safeguard privacy, balance efficiency with scalability, and develop trustworthy DataAgent guardrails to prevent malicious actions.
@@ 2. Background of Autonomous Data Agents
LLMs for Data Tasks. Recently, LLMs transforms this landscape. LLMs exhibit advanced reasoning and language comprehension, and can interact and adapt to external environments and verifier feedback. Researchers have generalized LLMs to translate ambiguous instructions from humans into concrete data operations and integrate with data tools to address feature selection (Gong et al., 2025a), feature transformation (Gong et al., 2025c), feature engineering (Abhyankar et al., 2025; Ko et al., 2025; Hollmann et al., 2023), data generation (Hsieh et al., 2025), texts to SQL and data retrieval (Liu et al., 2024), tabular data QA (Zhu et al., 2024c, a; Nguyen et al., 2024), data labeling (Wang et al., 2024a, a; Lu et al., 2025; Zhang et al., 2025), data cleaning (Biester et al., 2024; Li et al., 2024; Naeem et al., 2024; Zhang et al., 2024a), data preprocessing (Zhang et al., 2023; Meguellati et al., 2025). LLMs can reason and generate texts given a query in a reactive fashion via single-turn interactions, however, LLMs are not goal-driven systems in an autonomous fashion, and, thus, cannot plan, decompose, reason, and act to accomplish tasks.
@@ 4.1. Perception: Representing Task Descriptions and Data Contexts
title”: Albert Einstein”,
infobox”: {
Born”: ”14 March 1879”,
Died”: ”18 April 1955”,
Fields”: [”Theoretical physics”],
Known for”: [”Theory of relativity”, Photoelectric effect”]
text”: Albert Einstein was a German-born theoretical physicist…”,
images”: [
url”: https://upload.wikimedia.org/example.jpg”,
caption”: Einstein in 1921”
@@ 6.1. Primary Workflow
Overview. We evaluate a modular DataAgent that executes tabular ML tasks via a stepwise pipeline: (i) cleaning / preprocessing using a small rule-based controller (if–else rules over schema/profiled stats);  (ii) routing by a large LLM (e.g., GPT family) from a natural-language dataset card to one of three solution modes: classical linear/sparse model (e.g., Lasso), neural model (RL trained), or LLM-based sequence generator;  (iii) planning + tool calling (merged): for classical/neural, it calls a fixed tool chain to shape a better feature set; for LLM mode, it generates a sequence of operations (plan-as-sequence) and calls tools accordingly;  (iv) grounding that validates the called tools or the generated operation sequence (schema checks, dry-run, unit tests);  (v) execute to produce downstream predictions/plots;  (vi) evaluator that performs automatic sanity checks and triggers at most $K$ self-repair rounds. (vii) summary from the log of each step and generate a readable report for non-LLM-experts.
Figure 6 summarizes the overall workflow.
@@
6.2. Experimental Setup
Motivation. The experimental design is inspired by three key comparisons that highlight the role of modular agents in tabular ML tasks. First, by benchmarking against classical baselines, we measure whether the agent pipeline provides consistent performance gains. Second, by comparing to a pure LLM approach, we evaluate efficiency and automation: how many attempts are required to finish the task and how often grounding prevents silent failures. Third, by contrasting with the RL-based policies, we assess the benefits of being training-free while still retaining exploration diversity.
Datasets. We adopt regression tasks from OpenML and classification tasks from UCIrvine.
Methods. We evaluate: (i) DataAgent (ours): full workflow in Sec. 6.1; (ii) Baselines: fixed classical and neural models; (iii) Pure LLM: direct few-shot sequence generators without agent modules; (iv) RL-Policies: an RL-trained sequence explorer and an encoder-decoder policy.
Metrics. Primary metrics include: (1) downstream performance (1-RAE (Relative Absolute Error) for regression; F1-score for classification); (2) Task Success Rate (fraction of runs passing evaluator checks); (3) Attempts-to-Success (number of plans/calls needed until success); (4) Running Time to a valid solution.
6.3. Improvements
We present improvements of the proposed DataAgent from three perspectives: (i) Downstream performance, comparing predictive F1-score against classical, neural, LLM-only, and RL-based baselines across regression and classification tasks; (ii) Autonomous check, quantifying robustness via the number of attempts needed to complete a task and the rate of grounding errors; (iii) Efficiency check, reporting time and resource consumption versus alternatives. Together, these results highlight that the modular agent design not only yields higher task quality but also ensures automation and practical efficiency.
6.3.1. Downstream Performance
We first compare DataAgent with all baselines on regression and classification datasets (Table 3). Each row corresponds to one method, and each column is a dataset. The best performance for each dataset is highlighted in bold.
@@ 6.3.1. Downstream Performance
Source UCIrvine UCIrvine OpenML OpenML
Task Cls Cls Reg Reg
@@
NFS 68.67% 79.40% 0.5443 0.3473
AFAT 68.32% 76.03% 0.5435 0.2472
PCA 67.92% 70.92% 0.1109 0.1016
TTG 64.51% 76.03% 0.5443 0.3467
GRFG 68.29% 81.65% 0.5768 0.4562
FeatLLM 76.35% 80.07% 0.6477 0.4597
CAAFE 59.92% 70.60% N/A N/A
AutoFeat 74.86% 76.06% 0.6329 0.4407
OpenFE 74.50% 76.06% 0.6311 0.4402
ELLM-FT 76.39% 86.14% 0.6328 0.4734
@@
RL-Policy-1 (GRFG) 68.29% 81.65% 0.5768 0.4562
RL-Policy-2 (MOAT) 72.44% 86.95% 0.6251 0.4734
Ours (DataAgent) 79.60% 86.79% 0.7832 0.7488
Overall, Table 3 shows that DataAgent delivers the strongest downstream performance across both classification and regression benchmarks. On the German Credit dataset, it achieves 79.60%, exceeding all classical and LLM-based baselines. On SpectF, the performance is competitive with the best RL policy (MOAT), while still higher than most other baselines. For regression, DataAgent provides substantial gains: on OpenML 586 it reaches 0.7832, outperforming RL and Pure LLM variants, and on OpenML 618 it achieves 0.7488, the best among all compared methods. These results confirm that the agent workflow consistently yields improved predictive quality across different task types.
6.3.2. Autonomous Check
We evaluate autonomy by measuring the number of attempts required to reach a valid solution and the grounding error rate across datasets. As shown in Figure 7, DataAgent achieves the highest F1-score among all compared methods while requiring fewer attempts. In contrast, pure LLM baselines (Llama 3.1 8B and Llama 3 8B) often need multiple retries and still fail more frequently due to errors like non-executable sequences. These results highlight that the agent design not only improves predictive performance but also guarantees task completion within a small attempt budget, making it more reliable and user-friendly in practical data analysis scenarios.
@@ 6.3.2. Autonomous Check
Figure 7. Autonomous check: DataAgent achieves higher F1-score and lower grounding error rates than pure LLM baselines, ensuring reliable task completion.
6.3.3. Efficiency Check
We further evaluate training cost, inference latency, and the number of calls required to reach the best performance. As shown in Table 4, both RL policies incur substantial training overhead (hundreds to thousands of seconds), while our agent requires no training at all. During inference, DataAgent also operates with a much smaller number of calls (only 2 on average) compared to RL-based policies (13 and 22 calls, respectively). Despite its training-free nature and lightweight execution, DataAgent still achieves the best performance (0.7832), surpassing both RL-Policy-1 and RL-Policy-2. This demonstrates that the modular agent design leverages the knowledge embedded in pre-trained LLMs to handle diverse datasets more robustly.
@@ 6.3.3. Efficiency Check
Metric RL-Policy-1 (GRFG) RL-Policy-2 (MOAT) Ours (DataAgent)
@@
6.4. Insights
From the experimental results, we derive the following insights:
(1) Comparison with baselines.
Across both regression and classification tasks, DataAgent improves downstream performance over classical and AutoML-style baselines. The agent workflow ensures that the chosen modeling path better matches dataset characteristics and avoids failures. This demonstrates that the modular agent design is not only correct but also yields superior quality.
(2) Comparison with pure LLM.
Pure LLM approaches often require multiple attempts to arrive at a valid executable solution and suffer from grounding errors. By contrast, DataAgent integrates sequence grounding. This confirms the agent is end-to-end autonomous, guaranteeing completion, and making it usable by data specialists without LLM expertise.
(3) Comparison with RL-based agents.
While RL policies achieve competitive performance, they have a training cost for each dataset and require environment-specific feedback loops. DataAgent is training-free, adapting zero-shot to unseen datasets, while still enabling diverse exploration through LLM-based planning. This balance highlights that modular agent design can provide RL-level robustness without the heavy investment of policy training.
(4) Future work.
Experiments also suggest clear directions for improvement. First, the current routing module is unsupervised; supervised or weakly-supervised training could close the gap to oracle routing. Second, planning and tool-calling show that explicit plans reduce structural errors; future work could study plan granularity and toolbox scaling systematically. Third, evaluator logs provide the opportunity to build a memory module that reuses past repair trajectories to improve efficiency on similar future tasks.
@@ 8.1. Open Datasets and Benchmarks for Training DataAgents on Diverse Skills
Open datasets and benchmarks are essential for training data AI agents on task decompositions, action reasoning, action grounding, and training advanced skills like data preprocessing, data quality assessment, feature engineering, text-to-SQL, symbolic regression, tabular QA, and data repairs. There are benchmark datasets for specific skills: tabular QA (Zhong et al., 2017; Yu et al., 2018; Pasupat and Liang, 2015; Chen et al., 2019; Nan et al., 2022), text-to-SQL (Pasupat and Liang, 2015; Gan et al., 2021), symbolic equation extraction (Matsubara et al., 2022; de Franca et al., 2024). However, there is limited datasets and benchmarks for training agents to learn how to decompose data-related tasks, and other skills like feature engineering, data quality assessment, data repairs. The research community should curate diverse datasets reflecting real world data-related tasks.

### v2

@@ Abstract.
As data continues to grow in scale and complexity, preparing, transforming, and analyzing it remains labor-intensive, repetitive, and difficult to scale. Since data contains knowledge and AI learns knowledge from it, the alignment between AI and data is essential. However, data is often not structured in ways that are optimal for AI utilization. Moreover, an important question arises: how much knowledge can we pack into data through intensive data operations? Autonomous data agents (DataAgents), which integrate LLM reasoning with task decomposition, action reasoning and grounding, and tool calling, can autonomously interpret data task descriptions, decompose tasks into subtasks, reason over actions, ground actions into python code or tool calling, and execute operations. Unlike traditional data management and engineering tools, DataAgents dynamically plan workflows, call powerful tools, and adapt to diverse data tasks at scale. This report argues that DataAgents represent a paradigm shift toward autonomous data-to-knowledge systems. DataAgents are capable of handling collection, integration, preprocessing, selection, transformation, reweighing, augmentation, reprogramming, repairs, and retrieval. Through these capabilities, DataAgents transform complex and unstructured data into coherent and actionable knowledge. We first examine why the convergence of agentic AI and data-to-knowledge systems has emerged as a critical trend. We then define the concept of DataAgents and discuss their architectural design, training strategies, as well as the new skills and capabilities they enable. Finally, we call for concerted efforts to advance action workflow optimization, establish open datasets and benchmark ecosystems, safeguard privacy, balance efficiency with scalability, and develop trustworthy DataAgent guardrails to prevent malicious actions.
@@ 2. Background of Autonomous Data Agents
LLMs for Data Tasks. Recently, LLMs transforms this landscape. LLMs exhibit advanced reasoning and language comprehension, and can interact and adapt to external environments and verifier feedback. Researchers have generalized LLMs to translate ambiguous instructions from humans into concrete data operations and integrate with data tools to address feature selection (Gong et al., 2025a), feature transformation (Gong et al., 2025c), feature engineering (Abhyankar et al., 2025; Ko et al., 2025; Hollmann et al., 2023), data generation (Hsieh et al., 2025), texts to SQL and data retrieval (Liu et al., 2024), tabular data QA (Zhu et al., 2024c, a; Nguyen et al., 2024), data labeling (Wang et al., 2024a, a; Lu et al., 2025; Zhang et al., 2025), data cleaning (Biester et al., 2024; Li et al., 2024; Naeem et al., 2024; Zhang et al., 2024a), data preprocessing (Zhang et al., 2023a; Meguellati et al., 2025). LLMs can reason and generate texts given a query in a reactive fashion via single-turn interactions, however, LLMs are not goal-driven systems in an autonomous fashion, and, thus, cannot plan, decompose, reason, and act to accomplish tasks.
@@ 4.1. Perception: Representing Task Descriptions and Data Contexts
1{
2 "title": "Albert Einstein",
3 "infobox": {
4 "Born": "14 March 1879",
5 "Died": "18 April 1955",
6 "Fields": ["Theoretical physics"],
7 "Known for": ["Theory of relativity", "Photoelectric effect"]
8 },
9 "text": "Albert Einstein was a German-born theoretical physicist...",
10 "images": [
11 {
12 "url": "https://upload.wikimedia.org/example.jpg",
13 "caption": "Einstein in 1921"
14 }
15 ]
16}}
@@ 6.1. Primary Workflow
Overview. We evaluate a modular DataAgent that executes tabular ML tasks via a stepwise pipeline: (i) cleaning / preprocessing using a small rule-based controller (if–else rules over schema/profiled stats);  (ii) routing by a large LLM (e.g., GPT family) from a natural-language dataset card to one of three solution modes: classical linear/sparse model (e.g., Lasso (Tibshirani, 1996)), neural model (RL trained), or LLM-based sequence generator;  (iii) planning + tool calling (merged): for classical/neural, it calls a fixed tool chain to shape a better feature set; for LLM mode, it generates a sequence of operations (plan-as-sequence (Yao et al., 2023; Schick et al., 2023)) and calls tools accordingly;  (iv) grounding that validates the called tools or the generated operation sequence (schema checks, dry-run, unit tests);  (v) execute to produce downstream predictions/plots;  (vi) evaluator that performs automatic sanity checks and triggers at most $K$ self-repair rounds (Shinn et al., 2023; Madaan et al., 2023; Gu et al., 2024). (vii) summary from the log of each step and generate a readable report for non-LLM-experts. Figure 6 summarizes the overall workflow.
@@
6.2. Experimental Design: Comparing Paradigms
Purpose. Our goal is to use experiments as a lens to understand the strengths and weaknesses of three major paradigms for tabular data engineering: (1) classical and AutoML baselines, (2) pure LLM approaches, and (3) RL-based methods. DataAgent serves as a notable example of emerging modular agent paradigm, showing how agentic design balances reliability, autonomy, and efficiency.
Datasets. We include classification tasks from UCIrvine and regression tasks from OpenML, covering both discrete and continuous prediction settings.
Evaluation Aspects. We examine four aspects that highlight the trade-offs among paradigms: (i) Predictive quality, measured by 1-RAE for regression and F1-score for classification; (ii) Reliability, measured by task success rate and number of attempts needed; (iii) Automation, reflected in grounding error rates; (iv) Efficiency, measured by time and calls required.
6.3. Observations from Comparative Results
Predictive Quality Across Paradigms. Table 3 summarizes results across regression and classification datasets. Classical and AutoML baselines achieve solid but limited performance. Pure LLM approaches show strong results on some datasets, but suffer from instability across tasks. RL-based policies can reach high accuracy, yet often at the expense of extensive training. In contrast, DataAgent, as a modular agent instance, consistently delivers robust accuracy while avoiding training overhead, reflecting the potential of this paradigm for practical deployment in realistic scenarios.
Reliability and Automation. Figure 7 compares DataAgent with pure LLM baselines on task performance and error rates. Pure LLMs often require multiple retries to produce an executable solution and exhibit higher grounding error rates. In contrast, DataAgent achieves reliable task completion within a very small attempt budget (only 2 attempts on average), while maintaining the highest performance. This demonstrates how the modular agent design combines the reasoning ability of LLMs with structured grounding, ensuring both higher reliability and stronger automation.
Efficiency Trade-offs Table 4 compares only RL-based policies with the modular agent (DataAgent). RL-Policy-1 and RL-Policy-2 incur substantial training costs and require many calls during inference. DataAgent is training-free, reaches valid solutions with only 2 calls, and attains the strongest performance, despite a modest per-run inference latency. This shows that RL-based methods require heavy training, whereas the modular agent avoids training and remains efficient with only light inference overhead.
@@ 6.3. Observations from Comparative Results
Source UCIrvine (Public, 2023b) UCIrvine (Public, 2023b) OpenML (Public, 2023a) OpenML (Public, 2023a)
Task Classification Classification Regression Regression
@@
NFS (Chen et al., 2019a) 68.67% 79.40% 0.5443 0.3473
AFAT (Horn et al., 2020) 68.32% 76.03% 0.5435 0.2472
PCA (Maćkiewicz and Ratajczak, 1993) 67.92% 70.92% 0.1109 0.1016
TTG (Khurana et al., 2018) 64.51% 76.03% 0.5443 0.3467
FeatLLM (Han et al., 2024) 76.35% 80.07% 0.6477 0.4597
CAAFE (Hollmann et al., 2023) 59.92% 70.60% N/A N/A
AutoFeat (Horn et al., 2019) 74.86% 76.06% 0.6329 0.4407
OpenFE (Zhang et al., 2023b) 74.50% 76.06% 0.6311 0.4402
ELLM-FT (Gong et al., 2025b) 76.39% 86.14% 0.6328 0.4734
@@
RL-Policy-1 (GRFG (Wang et al., 2022)) 68.29% 81.65% 0.5768 0.4562
RL-Policy-2 (MOAT (Wang et al., 2023)) 72.44% 86.95% 0.6251 0.4734
DataAgent 79.60% 86.79% 0.7832 0.7488
@@
Figure 7. Comparison of modular agent (DataAgent) and pure LLM baselines, showing higher reliability and fewer retries for the modular agent paradigm.
@@
Metric RL-Policy-1 (GRFG (Wang et al., 2022)) RL-Policy-2 (MOAT (Wang et al., 2023)) DataAgent
@@
6.4. Takeaways on Paradigms for Tabular Data
From these comparisons, several insights emerge:
(1) Classical and AutoML baselines.
These methods are mature, stable, and widely used in practice. They provide a strong reference point, but their reliance on fixed model classes and handcrafted transformations limits adaptability to diverse and evolving datasets. As a result, they often plateau in performance when facing complex feature interactions.
(2) Pure LLM approaches.
LLMs show strong potential for automation by directly generating transformation or modeling sequences. However, their lack of grounding makes them prone to invalid outputs and silent failures, requiring multiple retries to succeed. This makes them less reliable as standalone solutions, though their flexibility remains valuable.
(3) RL-based policies.
RL methods can explore transformation spaces thoroughly and sometimes reach strong performance. Yet, they depend on heavy training for each dataset and require detailed environment-specific feedback loops. This investment in time and compute makes them less practical for broad deployment, despite their robustness.
(4) Modular agent paradigm (DataAgent).
The modular agent design strikes a balance across these trade-offs. It leverages LLMs for reasoning and exploration, but grounds outputs through explicit modules, ensuring task completion. Unlike RL, it is training-free and adapts zero-shot to new datasets, yet still retains exploratory diversity. This positions modular agents as a promising and practical paradigm for general-purpose tabular ML.
@@ 8.1. Open Datasets and Benchmarks for Training DataAgents on Diverse Skills
Open datasets and benchmarks are essential for training data AI agents on task decompositions, action reasoning, action grounding, and training advanced skills like data preprocessing, data quality assessment, feature engineering, text-to-SQL, symbolic regression, tabular QA, and data repairs. There are benchmark datasets for specific skills: tabular QA (Zhong et al., 2017; Yu et al., 2018; Pasupat and Liang, 2015; Chen et al., 2019b; Nan et al., 2022), text-to-SQL (Pasupat and Liang, 2015; Gan et al., 2021), symbolic equation extraction (Matsubara et al., 2022; de Franca et al., 2024). However, there is limited datasets and benchmarks for training agents to learn how to decompose data-related tasks, and other skills like feature engineering, data quality assessment, data repairs. The research community should curate diverse datasets reflecting real world data-related tasks.